OpenStudy (anonymous):

Jolie can read around 14 sentences per minute of her current novel. She wants to finish her reading before the movie comes out next week. If she has 17 pages left to read and each page has about 120 sentences, how many hours should Jolie set aside to complete her reading? Show your work PLEASE HELP

OpenStudy (anonymous):

Step 1: Find how many sentences there are in 17 pages. \[17 pages \times 120 sentences = 2040 sentences\]

OpenStudy (anonymous):

Step 2: Find how many minutes it takes to finish reading 2040 sentences. \[2040 sentences \div 14 sentences/minute \approx 145.7 minutes\]

OpenStudy (anonymous):

Step 3: Find how many hours are equivalent to 145.7 minutes. \[145.7 minutes \times \frac{ 1 hour }{ 60 minutes}\approx2.4 hours\]

OpenStudy (anonymous):

Or approximately 2 hours and 24 minutes.